For Trials 2 and 3, the percent CH_3COOH in vinega
Answer #1
Concepts and reason

The given problem is based on the concept of average percent and relative standard deviation (RSD) of a data.

Average of a data is equal to the sum of the data values divided by the total number of values in the data.

Relative standard deviation is a measure of precision with respect to the average value. It is expressed in percentage.

Fundamentals

The formula to calculate the average (xˉ)\left( {{\rm{\bar x}}} \right) is:

xˉ=x1+x2+x3.....xnn\bar x = \frac{{{x_1} + {x_2} + {x_3}.....{x_n}}}{n}

Here,

nn is the total number of value.

The formula to calculate standard deviation (σ)\left( \sigma \right) is:

σ=1n1Σ(xixˉ)2\sigma = \sqrt {\frac{1}{{n - 1}}\Sigma {{\left( {{x_{\rm{i}}} - \bar x} \right)}^2}}

The formula to calculate relative standard deviation (%RSD)\left( {{\rm{\% RSD}}} \right) is:

%RSD=(σxˉ)100%{\rm{\% RSD}} = \left( {\frac{\sigma }{{\bar x}}} \right)100{\rm{ \% }}

(5.a)

The formula to calculate the average percent is:

xˉ=x1+x22\bar x = \frac{{{x_1} + {x_2}}}{2} …… (1)

Substitute 5.01%5.01{\rm{ \% }} for x1{x_1} and 4.66%{\rm{4}}{\rm{.66 \% }} for x2{x_2} in equation (1).

xˉ=(5.01%)+(4.66%)2=4.84%\begin{array}{c}\\\bar x = \frac{{\left( {5.01{\rm{ \% }}} \right) + \left( {{\rm{4}}{\rm{.66 \% }}} \right)}}{2}\\\\ = 4.84{\rm{ \% }}\\\end{array}

(5.b)

The formula to calculate standard deviation (σ)\left( \sigma \right) is:

σ=(x1xˉ)2+(x2xˉ)2n1\sigma = \sqrt {\frac{{{{\left( {{x_{\rm{1}}} - \bar x} \right)}^2} + {{\left( {{x_{\rm{2}}} - \bar x} \right)}^2}}}{{n - 1}}} …… (2)

Substitute 5.01%5.01{\rm{ \% }} for x1{x_1} , 4.66%{\rm{4}}{\rm{.66 \% }} for x2{x_2} , 4.84%4.84{\rm{ \% }} for xˉ\bar x and 2 for nn in equation (2).

σ=(5.01%4.84%)2+(4.66%4.84%)221=(0.0289%)+(0.0324%)1=0.248%\begin{array}{c}\\\sigma = \sqrt {\frac{{{{\left( {5.01{\rm{ \% }} - 4.84{\rm{ \% }}} \right)}^2} + {{\left( {{\rm{4}}{\rm{.66 \% }} - 4.84{\rm{ \% }}} \right)}^2}}}{{2 - 1}}} \\\\ = \sqrt {\frac{{\left( {0.0289{\rm{ \% }}} \right) + \left( {0.0324{\rm{ \% }}} \right)}}{1}} \\\\ = 0.248{\rm{ \% }}\\\end{array}

The formula to calculate relative standard deviation (%RSD)\left( {{\rm{\% RSD}}} \right) is:

%RSD=(σxˉ)100%{\rm{\% RSD}} = \left( {\frac{\sigma }{{\bar x}}} \right)100{\rm{ \% }} …… (3)

Substitute 0.248%0.248{\rm{ \% }} for σ\sigma and 4.84%4.84{\rm{ \% }} for xˉ\bar x in equation (3).

%RSD=(0.248%4.84%)100%=5.12%\begin{array}{c}\\{\rm{\% RSD}} = \left( {\frac{{0.248{\rm{ \% }}}}{{4.84{\rm{ \% }}}}} \right)100{\rm{ \% }}\\\\ = {\rm{5}}{\rm{.12 \% }}\\\end{array}

Ans: Part 5.a

The average percent of acetic acid (CH3COOH)\left( {{\bf{C}}{{\bf{H}}_{\bf{3}}}{\bf{COOH}}} \right) in the vinegar sample is 4.84%{\bf{4}}.{\bf{84 \% }} .